Suppose that you and a friend are playing cards and decide to make a bet. If your friend draws three diamonds in succession from a standard deck of 52 cards with replacement, you give him $50 . Otherwise, he pays you $10. If the same bet was made 20 times, how much would you expect to win or lose? Round your answer to the nearest cent, if necessary.
Answer #1

The probability of the three diamonds

=13/52*12/51*11/50=0.0129

P(winning) = 0.9871

Expected value in one term = 0.0129*-50 + 0.9871*10

=-0.645+9.8715

=9.2256

In 20 terms the profit will be . = 20*9.2256=184.512